How do you find out the perimeter of a right triangle if you are only given the longest and shortest side?
The longest side of my triangle is 25in

The shortest side is 15in

And the medium side is unknow

Answer by dfrazzetto(277)


It's a 3, 4, 5 right triangle, with every side scaled (multiplied) by 5
medium side is 4*5 = 20
CHECK via pythagorean theorem:
a^2 + b^2 = c^2
15^2 + 20^2 = 25^2
225 + 400 = 625